a ladder is leaning against a house. the bottom of the ladder is 8 feet from the house. the top of the ladder is leaning against the house 15 feet above the ground. find the slope of the ladder.
please solve this word problem!

Answer by rothauserc(4718)

You have two points given, (8,0) and (0,15) the two points correspond to the distance from the base of the building to the bottom of the ladder and the vertical height to the top of the ladder leaning against the building.
***************************************************
we know that slope = (y2 - y1) / (x2 -x1) = (15 - 0) / (0 - 8) = -1.875